OpenStudy (anonymous):

• When should you solve a system of equations by graphing? • When should you solve a system of equations by elimination? • When should you solve a system of equations by substitution?

OpenStudy (campbell_st):

If you are looking at linear equations it really its at your choice... you can solve any system of equations by any of the methods... quadratic or higher and a linear, I use substitution most of the time... but again, this can be done graphically you can solve any equation graphically e.g. x + 3 = 5 just graph y = x + 3 and y = 5... it will show you x = 2

OpenStudy (anonymous):

Depends on what the question looks like. Elimination is usually the easiest, but it gets tricky sometimes. I only use graphing as a last option and professor dont usually ask this

OpenStudy (anonymous):

Those wt really know how to answer them. asked me , i dont really know how to answer them each individually

hero (hero):

If you're given: `x` = 3 + y 2x + 5y = 7 Use `substitution` since in this case, x is isolated and you can substitute the expression for x in the first equation into the second equation. However if you're given: x `+ 2y` = 4 3x `- 2y` = 5 Use `elimination` since you can simply add both equations together to get rid of the y variable.

hero (hero):

If given `y` = 2x - 3 `y` = 5x - 4 Then use `graphing` since both equations are in the form y = mx + b. You can simply enter both equations into your graphing calculator to find the (x,y) intersection point.

OpenStudy (anonymous):

Thank you.